OpenStudy (anonymous):

Find the x and y intercepts. Then graph the equation. 25y-5x=-10

OpenStudy (anonymous):

X-INTERCEPT, Y=0, 0-5x=-10 x=2 (2,0) Y-INTERCEPT, X=0 25y=-10 y=-0.4 (0,-0.4)